# System-Embedded Diffusion Bridge Models

### v1

@@ 6 Discussion and limitations
[…]
Appendix A Proofs and derivations
@@ A.1 Derivation of drift and diffusion coefficients for SDB
We begin with deriving the drift and diffusion coefficients $\mathbf{F}_{t},\mathbf{G}_{t}$ of SDB (eqs. 10a and 10b) by using theorem 1 with the mean and covariance matrices $\mathbf{H}_{t},\boldsymbol{\Sigma}_{t}$ (eqs. 8a and 8b) of its forward process. Recall that $\mathbf{H}_{t}=\mathbf{A}^{+}\mathbf{A}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A}),\boldsymbol{\Sigma}_{t}=\gamma_{t}\mathbf{A}^{+}\boldsymbol{\Sigma}{\mathbf{A}^{+}}^{\top}+\beta_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$. Following theorem 1, we obtain
@@
Moreover, note that $\mathbf{F}_{t}^{\top}=\mathbf{F}_{t},\boldsymbol{\Sigma}_{t}^{\top}=\boldsymbol{\Sigma}_{t}$ due to the symmetry of the range and null space projections, and $\mathbf{F}_{t}\boldsymbol{\Sigma}_{t}=\boldsymbol{\Sigma}_{t}\mathbf{F}_{t}$ since $\mathbf{F}_{t}$ only affects the null space and $(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})^{2}=(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$ (idempotent). Following theorem 1,
@@ A.2 Proof of theorem 2
(Analytic posterior given boundary pair) [Liu et al., 2023] The posterior of eq. 1 given some boundary pair $(\mathbf{x}_{0},\mathbf{x}_{1})$ admits an analytic form:
@@
where $\mathbf{v}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\frac{g^{2}(t)}{\sigma_{t}^{2}}(\mathbf{x}_{t}-\mathbf{x}_{0})$ whose solution is the posterior mean of eq. 14.
[…]
@@
Equivalently, $(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{t}\sim\boldsymbol{\mathcal{N}}((\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\left(\frac{\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}\mathbf{x}_{0}+\frac{\sigma_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}\mathbf{x}_{1}\right),\frac{\sigma_{t}^{2}\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A}))$. Hence, the posterior mean of SDB (SB) takes the form stated in proposition 1 and proposition 2 can be applied directly to the null space part.
As a final remark, we note that, up to this point, the definition of $g(t)$ is interdependent with that of $\alpha_{t}$ and $\beta_{t}$. It is not immediately clear that defining $\alpha_{t},\beta_{t}$ as in theorem 2 fulfills $g^{2}(t)=\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}$. We now show that this property is indeed satisfied.
@@ B.1 Diffusion models for inverse problems
In recent years, deep neural networks have gained significant attention for solving various inverse problems from different perspectives [Ongie et al., 2020, Chen et al., 2022]. Solving eq. 4 can be interpreted as sampling from the posterior distribution $p(\mathbf{x}\mid\mathbf{y})$, and generative models that support conditional generation are naturally suited to this task. Diffusion models, in particular, have emerged as SOTA tools for this purpose, thanks to their flexible mathematical formulation and expressive data priors [Daras et al., 2024]. The standard approach involves extending the diffusion process to sample from $p(\mathbf{x}_{0}\mid\mathbf{y})$ at $t=0$ by applying Bayes’ Theorem. The conditional score function can be decomposed as $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t}\mid\mathbf{y})}}=\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}+\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{y}\mid\mathbf{x}_{t})}}$, where the first term, $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$, can be approximated using a pretrained score network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$, and the second term models the relationship between $\mathbf{x}_{t}$ and the measurement $\mathbf{y}$ [Song et al., 2021b, Dhariwal and Nichol, 2021]. This framework has spurred considerable progress in the field, with numerous successful methods emerging [Kawar et al., 2022, Chung et al., 2022, 2023a, Song et al., 2023a, Mardani et al., 2024, Chung et al., 2024]. Since this approach does not require retraining the score network for each specific problem, we refer to these methods as unsupervised.
@@
In a related approach, Luo et al. [2023a] derive a scalar case of eq. 1 that incorporates the start and end points, termed the mean-reverting SDE (IR-SDE). Through specific parameterization, they show that its score function is analytically tractable, connecting to the Ornstein-Uhlenbeck (OU) process [Gillespie, 1996]. Yue et al. [2024] extend this work with a generalized OU process and Doob’s h-transform (GOUB). Further extensions include Luo et al. [2023b], who apply IR-SDE in latent spaces, and Welker et al. [2022] and Richter et al. [2023], who adapt similar processes for speech tasks. Recently, Zhu et al. [2025] unify GOUB and DDBM within the framework of stochastic optimal control.
@@ B.3.1 Unsupervised
Pseudoinverse-Guided Diffusion Model ($\Pi$GDM, Song et al. [2023a]) proposes to approximate the loglikelihood score with $(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0}(\mathbf{x}_{t}))^{\top}(r_{t}^{2}\mathbf{A}\mathbf{A}^{\top}+\sigma_{\mathbf{y}}^{2}\mathbf{I})^{-1}\mathbf{A}\frac{\partial\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}$, where $\sigma_{\mathbf{y}}^{2}\mathbf{I}$ is the measurement system covariance and $r_{t}^{2}$ is a time-dependent term that should depend on the data (Appendix A.3., Song et al. [2023a]). Using automatic differentiation, one can compute $\frac{\partial\hat{\mathbf{x}}_{0}(\mathbf{x}_{t})}{\partial\mathbf{x}_{t}}$. By combining it with the other terms, the entire approximation can be efficiently computed as a vector-Jacobian product.
@@ B.3.2 Supervised
[…]
@@
[…]
@@ C.2 Nonlinear measurement systems
While not the focus of this work, inverse problems with nonlinear system response $\mathcal{\mathbf{A}}$ may also be considered, e.g., in the problem of GAN inversion [Creswell and Bharath, 2018]. Assuming that $\mathcal{\mathbf{A}}$ is differentiable, it may be locally linearized using Taylor expansion as $\mathcal{\mathbf{A}}(\mathbf{x})\approx\mathcal{\mathbf{A}}(\mathbf{x}_{0})+\mathcal{\boldsymbol{J}}_{\mathcal{\mathbf{A}}}\mathbf{x}_{0}(\mathbf{x}-\mathbf{x}_{0})$ around some point $\mathbf{x}_{0}$, where $\mathcal{\boldsymbol{J}}_{\mathcal{\mathbf{A}}}$ is the Jacobian of $\mathcal{\mathbf{A}}$. To adapt SDB to such cases, one can use $\mathcal{\boldsymbol{J}}_{\mathbf{A}}$ as a local linear approximation of $\mathbf{A}$ and treat it instead as the system response matrix. We consider further study of such models as future work.
C.3 Pseudocode
We provide pseudocode for SDB for both training and Euler-Maruyama sampling in algorithm 1 and algorithm 2 respectively.
Algorithm 1 SDB Training
1:$p(\mathbf{x}_{0}),p(t),\mathbf{A},\mathbf{A}^{+},\boldsymbol{\Sigma}^{1/2},\alpha_{t},\beta_{t},\gamma_{t},\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}}$
2:for each iteration do
3:     $\mathbf{x}\sim p(\mathbf{x})$
4:     $t\sim p(t)$
5:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{m}$
6:     $\boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{d}$
7:     $\mathbf{x}_{t}\leftarrow[\mathbf{A}^{+}\mathbf{A}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})]\mathbf{x}+\gamma_{t}^{\frac{1}{2}}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon}+\beta_{t}^{\frac{1}{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime}$
8:     $L_{\theta}\leftarrow\left\|\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}\right\|_{1}$
9:     $\theta\leftarrow\text{optimizer}(\nabla_{\theta}L_{\theta})$
10:end for
11:return $\theta$
Algorithm 2 SDB Sampling (Euler-Maruyama)
1:$N,\mathbf{A},\mathbf{A}^{+},\boldsymbol{\Sigma}^{1/2},\alpha_{t},\beta_{t},\gamma_{t},\mathbf{H}_{t},\mathbf{F}_{t},\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}},\mathbf{y}$
2:$t\leftarrow 1$
3:$\Delta t\leftarrow 1/N$
4:$\boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{d}$
5:$\hat{\mathbf{x}}\leftarrow\mathbf{A}^{+}\mathbf{y}$
6:$\mathbf{x}_{t}\leftarrow\hat{\mathbf{x}}+\beta_{t}^{\frac{1}{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime}$
7:for $i\in\{1,\ldots,N\}$ do
8:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{m}$
9:     $\boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\in\mathbb{R}^{d}$
10:      $\mathbf{x}_{t-\Delta t}\leftarrow\mathbf{x}_{t}+\Delta t\left[(f_{t}\mathbf{I}-2\mathbf{F}_{t})(\mathbf{H}_{t}\mathcal{\boldsymbol{D}}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)-\mathbf{x}_{t})-\mathbf{F}_{t}\mathbf{x}_{t}\right]+\\ \text{\hskip 40.00006pt}\Delta t^{1/2}\left[\left(\frac{\text{d}\gamma_{t}}{\text{d}t}\right)^{\frac{1}{2}}\mathbf{A}^{+}\boldsymbol{\Sigma}^{\frac{1}{2}}\boldsymbol{\epsilon}+\left(\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}\right)^{\frac{1}{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\boldsymbol{\epsilon}^{\prime}\right]$
11:     $t\leftarrow t-\Delta t$
12:end for
13:return $\mathbf{x}_{0}$

### v2

@@ 80. Guidelines:
Appendix A Theoretical results, proofs and derivations
@@ A.1 Derivation of drift and diffusion coefficients for SDB
We begin with deriving the drift and diffusion coefficients $\mathbf{F}_{t},\mathbf{G}_{t}$ of SDB (eqs.˜10a and 10b) by using theorem˜1 with the mean and covariance matrices $\mathbf{H}_{t},\boldsymbol{\Sigma}_{t}$ (eqs.˜8a and 8b) of its forward process. Recall that $\mathbf{H}_{t}=\mathbf{A}^{+}\mathbf{A}+\alpha_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A}),\boldsymbol{\Sigma}_{t}=\gamma_{t}\mathbf{A}^{+}\boldsymbol{\Sigma}{\mathbf{A}^{+}}^{\top}+\beta_{t}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$. Following theorem˜1, we obtain
@@
Moreover, note that $\mathbf{F}_{t}^{\top}=\mathbf{F}_{t},\boldsymbol{\Sigma}_{t}^{\top}=\boldsymbol{\Sigma}_{t}$ due to the symmetry of the range and null space projections, and $\mathbf{F}_{t}\boldsymbol{\Sigma}_{t}=\boldsymbol{\Sigma}_{t}\mathbf{F}_{t}$ since $\mathbf{F}_{t}$ only affects the null space and $(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})^{2}=(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$ (idempotent). Following theorem˜1,
@@ A.2 Proof of theorem˜2
(Analytic posterior given boundary pair) [Liu et al., 2023] The posterior of eq.˜1 given some boundary pair $(\mathbf{x}_{0},\mathbf{x}_{1})$ admits an analytic form:
@@
where $\mathbf{v}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\frac{g^{2}(t)}{\sigma_{t}^{2}}(\mathbf{x}_{t}-\mathbf{x}_{0})$ whose solution is the posterior mean of eq.˜14.
Consider the null space part of $\mathbf{x}_{t}$ given by mean and covariance matrices from eqs.˜8a and 8b with $\alpha_{t}=\frac{\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}},\beta_{t}=\frac{\sigma_{t}^{2}\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}$ for $\sigma_{t}^{2}=\int_{0}^{t}g^{2}(\tau)d\tau,\bar{\sigma}_{t}^{2}=\int_{t}^{1}g^{2}(\tau)d\tau$, where $g(t)$ is the null space diffusion coefficient given by eq.˜10b, i.e., $g(t)=\left(\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}\right)^{\frac{1}{2}}$:
@@
Equivalently, $(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\mathbf{x}_{t}\sim\boldsymbol{\mathcal{N}}((\mathbf{I}-\mathbf{A}^{+}\mathbf{A})\left(\frac{\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}\mathbf{x}_{0}+\frac{\sigma_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}\mathbf{x}_{1}\right),\frac{\sigma_{t}^{2}\bar{\sigma}_{t}^{2}}{\bar{\sigma}_{t}^{2}+\sigma_{t}^{2}}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A}))$. Hence, the posterior mean of SDB (SB) takes the form stated in proposition˜1 and proposition˜2 can be applied directly to the null space part.
As a final remark, we note that, up to this point, the definition of $g(t)$ is interdependent with that of $\alpha_{t}$ and $\beta_{t}$. It is not immediately clear that defining $\alpha_{t},\beta_{t}$ as in theorem˜2 fulfills $g^{2}(t)=\frac{\text{d}\beta_{t}}{\text{d}t}-2\beta_{t}\frac{\text{d}}{\text{d}t}\log{\alpha_{t}}$. We now show that this property is indeed satisfied.
@@
A.3 Proof of theorem˜3
Many diffusion bridge methods, such as I2SB, GOUB, or DDBM, rely on non-Markovian stochastic processes, where $\mathbf{x}_{t}$ depends on both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$. In contrast, SDB uses a Markovian forward diffusion process, for which $p(\mathbf{x}_{t}|\mathbf{x}_{0})$ is well-defined. More specifically, SDB is a special case of score-based generative models, using the general formulation with matrix-valued drift and diffusion coefficients described in Appendix A of [Song et al., 2021b]. We show below that SDB is a principled posterior sampler of $p(\mathbf{x}|\mathbf{y})$ with a proof of theorem˜3 based on a series of lemmas.
We start by considering the most general practical case with a positive semidefinite covariance matrix $\boldsymbol{\Sigma}$ in the forward model $p(\mathbf{y}^{\prime}\mid\mathbf{x})=\mathcal{N}(\mathbf{A}^{\prime}\mathbf{x},\boldsymbol{\Sigma})$. When $\boldsymbol{\Sigma}$ is positive definite, we can perform whitening by scaling the measurements and the system matrix as $\mathbf{y}=\boldsymbol{\Sigma}^{-\frac{1}{2}}\mathbf{y}^{\prime}$ and $\mathbf{A}=\boldsymbol{\Sigma}^{-\frac{1}{2}}\mathbf{A}^{\prime}$, yielding the equivalent isotropic model $p(\mathbf{y}\mid\mathbf{x})=\mathcal{N}(\mathbf{A}\mathbf{x},\mathbf{I})$. In the singular case (i.e., $\boldsymbol{\Sigma}$ is positive semidefinite but not positive definite), $\boldsymbol{\Sigma}^{-\frac{1}{2}}$ can be replaced with $(\boldsymbol{\Sigma}^{+})^{\frac{1}{2}}$ to achieve whitening on the range space of $\boldsymbol{\Sigma}$. From this point onward, we drop primes for notational simplicity and proceed in the whitened coordinates. Hence, without loss of generality, we assume an isotropic noise model, which is equivalent to an initial whitening of both the measurements and the system response matrix.
Lemma 1
The pseudoinverse reconstruction $\mathbf{A}^{+}\mathbf{y}$ is a sufficient statistic for estimation of $\mathbf{x}$ given $\mathbf{y}$. That is,
$p(\mathbf{x}|\mathbf{y})=p(\mathbf{x}|\mathbf{A}^{+}\mathbf{y}).$
This holds because (i) the range space component of the measurements can be exactly recovered by re-applying $\mathbf{A}$ to $\mathbf{A}^{+}\mathbf{y}$, and (ii) the null space component contains no useful information about $\mathbf{x}$.
Lemma 2
A sample from
$p(\mathbf{z}|\mathbf{y})=\mathcal{N}(\mathbf{A}^{+}\mathbf{y},\beta_{1}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A}))$
is also a sufficient statistic for estimating $\mathbf{x}$ given $\mathbf{y}$, i.e.,
$p(\mathbf{x}|\mathbf{y})=p(\mathbf{x}|\mathbf{z}).$
The null-space component of the pseudoinverse reconstruction is zero; therefore, adding null-space noise preserves sufficiency.
Lemma 3
The SDB forward process at the final time step, $p(\mathbf{x}_{1}|\mathbf{x}_{0}=\mathbf{x})$, is identically distributed to the SDB initializer $p(\mathbf{z}|\mathbf{x})$ from lemma˜2, assuming
$\lim_{t\rightarrow 1}\gamma_{t}=1,\quad\lim_{t\rightarrow 1}\frac{\alpha_{t}^{2}}{\beta_{t}}=0.$
Both are conditional Gaussian random vectors with mean $\mathbf{A}^{+}\mathbf{A}\mathbf{x}$ and covariance $\mathbf{A}^{+}\Sigma{\mathbf{A}^{+}}^{\top}+\beta_{1}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})$.
Lemma 4
The trained SDB model is a principled sampler of $p(\mathbf{x}_{0}|\mathbf{x}_{1})$.
Since SDB is a subset of score-based generative models with a standard Markovian forward process, the reverse SDE with a learned score network asymptotically samples from $p(\mathbf{x}_{0}|\mathbf{x}_{1})$ [Anderson, 1982], approaching the exact score with sufficient data and model capacity.
The procedure to generate principled posterior samples from $p(\mathbf{x}|\mathbf{y})$ is as follows:
  1. 1.
    Compute the pseudoinverse reconstruction $\mathbf{A}^{+}\mathbf{y}$.
  2. 2.
    Sample null-space noise:
    $\mathbf{z}\sim p(\mathbf{z}|\mathbf{y})=\mathcal{N}(\mathbf{A}^{+}\mathbf{y},\beta_{1}(\mathbf{I}-\mathbf{A}^{+}\mathbf{A})).$
  3. 3.
    Sample the SDB reverse process:
    $\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{1}=\mathbf{z}).$
By lemmas˜1 and 2, $\mathbf{z}$ is a sufficient statistic for $\mathbf{x}$. Lemma˜3 ensures that initializing the reverse process at $\mathbf{x}_{1}=\mathbf{z}$ is valid. Lemma˜4 guarantees that the trained SDB model samples from $p(\mathbf{x}_{0}|\mathbf{x}_{1})$. Combining these results, the procedure produces principled posterior samples from $p(\mathbf{x}|\mathbf{y})$.
$\square$
A.4 Magnus expansion generality
The matrix integral formulation of the reverse-time SDB process holds under the condition that the drift matrices $\mathbf{F}_{t}$ commute for all $t\in[0,1]$. In the proposed SDB formulation (parameterized as in eq.˜10a), this condition is always satisfied: the only time-dependent parameter is the scalar $\alpha_{t}$, so the singular vectors of $\mathbf{F}_{t}$ remain constant, ensuring commutativity at all times.
This property extends to all scalar diffusion processes, including VP, VE, Flow Matching [Lipman et al., 2023], Fourier Diffusion Models Tivnan et al. [2025], and Subspace Diffusion Models Jing et al. [2022]. While diffusion methods with non-commuting $F_{t}$ matrices are currently not known to us, theorem˜1 remains valid in those cases through the Magnus expansion Kamm et al. [2021].
@@ B.1 Diffusion models for inverse problems
In recent years, deep neural networks have gained significant attention for solving various inverse problems from different perspectives [Ongie et al., 2020, Chen et al., 2022]. Solving eq.˜4 can be interpreted as sampling from the posterior distribution $p(\mathbf{x}\mid\mathbf{y})$, and generative models that support conditional generation are naturally suited to this task. Diffusion models, in particular, have emerged as SOTA tools for this purpose, thanks to their flexible mathematical formulation and expressive data priors [Daras et al., 2024, Preechakul et al., 2022, Sobieski and Biecek, 2024]. The standard approach involves extending the diffusion process to sample from $p(\mathbf{x}_{0}\mid\mathbf{y})$ at $t=0$ by applying Bayes’ Theorem. The conditional score function can be decomposed as $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t}\mid\mathbf{y})}}=\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}+\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{y}\mid\mathbf{x}_{t})}}$, where the first term, $\nabla_{\mathbf{x}_{t}}{\log{p(\mathbf{x}_{t})}}$, can be approximated using a pretrained score network $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$, and the second term models the relationship between $\mathbf{x}_{t}$ and the measurement $\mathbf{y}$ [Song et al., 2021b, Dhariwal and Nichol, 2021]. This framework has spurred considerable progress in the field, with numerous successful methods emerging [Kawar et al., 2022, Chung et al., 2022, 2023a, Song et al., 2023a, Mardani et al., 2024, Chung et al., 2024, Sobieski et al., 2025]. Since this approach does not require retraining the score network for each specific problem, we refer to these methods as unsupervised.
@@
In a related approach, Luo et al. [2023a] derive a scalar case of eq.˜1 that incorporates the start and end points, termed the mean-reverting SDE (IR-SDE). Through specific parameterization, they show that its score function is analytically tractable, connecting to the Ornstein-Uhlenbeck (OU) process [Gillespie, 1996]. Yue et al. [2024] extend this work with a generalized OU process and Doob’s h-transform (GOUB). Further extensions include Luo et al. [2023b], who apply IR-SDE in latent spaces, and Welker et al. [2022] and Richter et al. [2023], who adapt similar processes for speech tasks. Recently, Zhu et al. [2025] unify GOUB and DDBM within the framework of stochastic optimal control.
@@ B.3.1 Unsupervised
[…]
@@ B.3.2 Supervised
where $\theta_{t}=\frac{\sigma^{2}_{t}}{\lambda^{2}},\bar{\theta}_{t}=\int_{0}^{t}{\theta_{\tau}}d\tau$ and $\lambda$ is a predefined constant. At $t=1$, IR-SDE arrives at a Gaussian distribution (with non-zero covariance) centered at $\mathbf{x}_{T}$.
@@
where $\theta_{t}=\frac{g^{2}(t)}{2\lambda^{2}},\bar{\theta}_{s:t}=\int_{s}^{t}\theta_{\tau}d\tau,\bar{\sigma}_{s:t}^{2}=\frac{g^{2}(t)}{2\theta_{t}}(1-e^{-2\bar{\theta}_{s:t}})$ for a predefined constant $\lambda$.
@@ C.2 Nonlinear measurement systems
While SDB is primarily formulated for linear operators $\mathbf{A}$ with a global range-nullspace decomposition, it can be extended to certain nonlinear inverse problems. For a nonlinear, differentiable system response $\mathcal{\mathbf{A}}$, we can approximate it locally using a first-order Taylor expansion around some point $\mathbf{x}_{0}$:
$\mathcal{\mathbf{A}}(\mathbf{x})\approx\mathcal{\mathbf{A}}(\mathbf{x}_{0})+\mathcal{\boldsymbol{J}}_{\mathcal{\mathbf{A}}}(\mathbf{x}_{0})(\mathbf{x}-\mathbf{x}_{0}),$ (26)
where $\mathcal{\boldsymbol{J}}_{\mathcal{\mathbf{A}}}(\mathbf{x}_{0})$ is the Jacobian of $\mathcal{\mathbf{A}}$ at $\mathbf{x}_{0}$. This linearization induces a local range-nullspace decomposition, enabling the use of SDB on nonlinear systems. The procedure is as follows:
  1. 1.
    For a measurement $\mathbf{y}=\mathcal{\mathbf{A}}(\mathbf{x})+\boldsymbol{\Sigma}^{1/2}\epsilon$, compute an approximate maximum likelihood estimate $\hat{\mathbf{x}}$ of the signal $\mathbf{x}$.
  2. 2.
[…]
  3. 3.
    Treat $\mathbf{y}\approx\mathbf{J}_{\mathbf{A}}(\hat{\mathbf{x}})\mathbf{x}+\boldsymbol{\Sigma}^{1/2}\epsilon$ as a linear Gaussian model and apply standard SDB with pseudoinverse and range/nullspace projections.
This approach leverages the Jacobian as a local linear approximation of the nonlinear operator, effectively adapting SDB to more general, differentiable systems.